# About the Burton-Miller factor in the low frequency region

### v1

@@ Abstract
The Burton-Miller method is a widely used approach in acoustics to enhance the stability of the boundary element method for exterior Helmholtz problems at so-called critical frequencies. This method depends on a coupling parameter $\eta$ and it can be shown that as long as $\eta$ has an imaginary part different from 0, the boundary integral formulation for the Helmholtz equation has a unique solution at all frequencies. A popular choice for this parameter is $\eta=\frac{{\mathrm{i}}}{k}$, where $k$ is the wavenumber. It can be shown that this choice is quasi optimal. However, especially in the low frequency region, where the critical frequencies are still sparsely distributed, different choices for this factor result in a smaller condition number and a smaller error of the solution. In this work, alternative choices for this factor are compared based on numerical experiments. Additionally, a way to enhance the Burton-Miller solution with $\eta=\frac{{\mathrm{i}}}{k}$ for a sound hard scatterer in the low frequency region by an additional step of a modified Richardson iteration is introduced.
@@ 1 Introduction
The boundary element method (BEM, [15, 16]) is an often used numerical method to solve exterior scattering problems in acoustics. It is a well known fact that although the exterior Helmholtz equation has a unique solution, there are certain fictitious frequencies where the boundary integral equation (BIE) used by the BEM has not [6, 17]. As the number of these critical frequencies below a given wavenumber $k$ is proportional to $k^{3}$ (cf. [2]), this non-uniqueness poses a major problem for the Helmholtz BEM.
@@
$\phi({\bf x})$ denotes the velocity potential at a point ${\bf x}$ on a smooth part of the boundary $\Gamma$, $v({\bf x}):=\frac{\partial\phi}{\partial{\bf n}_{x}}\left({\bf x}\right):=\nabla\phi({\bf x})\cdot{\bf n}_{x}$ and
@@
where ${\bf n}_{y}$ is the normal vector to the boundary $\Gamma$ at a point ${\bf y}\in\Gamma$.
@@
Different values for the BM factor $\eta$ and its sign have been discussed in the past (cf. [1, 2, 11, 14, 19, 21]), the most commonly used factor is $\eta=\frac{{\mathrm{i}}}{k}$, which is almost optimal in terms of having a small condition number for problems on the unit circle or the unit sphere. Since scattering problems involving the sphere have an analytic solution based on spherical harmonics and the spherical Bessel and Hankel functions, it is possible to derive the eigenvalues and the condition number of the linear system for the combined integral equation as a function of $\eta$. Based on this analysis, it can be shown that $\eta=\frac{{\mathrm{i}}}{k}$ is quasi-optimal (cf. [1]).
@@
$=\left(O\left(\frac{1}{r}\right)+O(k)\right)G({\bf x},{\bf y}),$
@@
$=\left(O\left(\frac{1}{r^{2}}\right)+O(k/r)+O(k^{2})\right)G({\bf x},{\bf y}),$
where $r=||{\bf x}-{\bf y}||$ (cf. [7, Eq. (15)]). The components of the dBIE contain a scaling with the wavenumber $k$, at least if $k$ is big enough so that the parts involving $O(k)$ and $O(k^{2})$ are bigger than the $O\!\left(\frac{1}{r^{2}}\right)$ and $O\!\left(\frac{1}{r}\right)$ parts. Using this viewpoint, multiplying the dBIE with $\frac{1}{k}$ balances the scale of BIE and dBIE. In 2D, a similar heuristic argument can be given when looking at the derivatives of the Green’s function $H_{0}(kr)$ with respect to $r$.
One can clearly see in Fig. 1 that for the example of plane wave scattering from a sound hard sphere, the Burton-Miller has a positive effect on the stability of the solution of the scattering problem. However, in the low frequency region, the relative error between numerical and analytical solution away from the critical frequencies is, in general, smaller for the version without Burton-Miller. Using again the heuristic viewpoint, it seems that in the low frequency region, the $O(k)$ factor becomes too small and the parts depending on the radius dominate the error.
The data in Fig. 1 was calculated for $N=1280$ almost regular triangles, using collocation with constant elements [12]. The highest investigated wavenumber of 500 Hz is slightly larger than suggested by the 6-to-8-elements per wavelength rule of thumb [13]. Still, the relative error at this frequency is acceptable.
@@
The example presented in Fig. 1 illustrates the problem of choosing a good BM-parameter if BEM solutions at low frequencies are needed. Examples where solutions for small wavenumbers are needed are the 2.5D BEM [9, 10] or the BEM for periodic structures using the Floquet transform [8]. The fact that for low frequencies the coupling factor $\eta=\frac{{\mathrm{i}}}{k}$ becomes very large, led to different approaches to define this factor for small $k$. One suggestion [1] is to use $\eta=\min(2,\frac{1}{k})\cdot{\mathrm{i}}$, which means for the unit sphere that above about 27 Hz the BM-factor is $\eta=\frac{{\mathrm{i}}}{k}$, below it is set constant. It was stated in [2] that “[…] 1 has often been used in practice”, where the authors referred to the imaginary part of the coupling factor. In [9], a factor $\eta=\frac{{\mathrm{i}}|k|}{\max(|k|^{2},50)}$ was suggested in connection with the 2D and 2.5D BEM. In [5], a BM factor of $\eta=\min(\frac{1}{3},\frac{\lambda}{D})\cdot{\mathrm{i}}=\min(\frac{1}{3},\frac{2\pi}{Dk})\cdot{\mathrm{i}}$ was used [^1] In fact, in [5] the authors scale the BIE, and the factor was adapted accordingly., where $D$ is the diameter of the scatterer. According to [5], this factor is based on optimizing the number of necessary GMRES iterations. [5] explicitly includes the diameter $D$ of the scatterer, whereas in some studies this scaling with the size of the object is implicitly assumed (cf. [1]: “If we denote $kR$ by $k^{*}$ and $\nu/R$ but $\nu^{*}$ then the problem is reduced to that of a unit sphere at the wave number $k^{*}$ with the coupling parameter $\nu^{*}$, a similar claim can be found in [11].).
@@
In this study, different choices of the Burton-Miller factor and their effect on the solution of a scattering problem will be investigated. As a benchmark example, the scattering of a plane wave on the unit sphere will be used. As there is some discussion about whether to adapt the BM factor to the size of the scatterer, one numerical experiment will look at the sphere with radius $r=2$. Also, a new approach based on a single step of a Modified Richardson iteration is proposed to enhance the solution of the Burton-Miller BEM in the low frequency region for sound hard problems.
2 Additional Iteration Step
@@
It was illustrated in the previous section that in the low frequency region away from the critical frequencies, the Burton-Miller method produces larger errors than the solution of the original BIE. An alternative to using different Burton-Miller factors is presented in the following. As the operators and integral kernels involved with the dBIE are, in general, numerically more challenging than the operators involved with the BIE, it is plausible to assume that the error of a solution based purely on the BIE is smaller than the error for the Burton-Miller solution as long as the frequency is sufficiently far away from the set of critical frequencies, where the BIE has no unique solution (see also Fig. 1).
@@
In [18, Theorem 3] it is shown that if a system $\phi-K\phi=0$ with compact operator $K$ has only the trivial solution $\phi=0$, and if $\phi_{0}$ is a Galerkin-Petrov solution of $\phi=f+K\phi$, then $\phi_{1}=f+K\phi_{0}$ is a better approximation of the real solution provided that the number of ansatz functions $n$ is big enough. As the theorem is based on orthogonal projections on the ansatz and test function spaces, it cannot be applied to the collocation method directly. Also, the fact that the system needs to have a unique solution is not fulfilled at the critical frequencies. Nevertheless, this theorem serves as a motivation for using an additional iteration step in connection with the problem of solving wave scattering around a sound hard sphere:
@@ 3.1 Sound Hard Sphere
As the first benchmark example, the standard problem of plane wave scattered at a sound hard sphere is used. This example has the advantage, that an analytic solution of the problem exists [20, Eq. (6.186)], and it is therefore possible to easily calculate the numerical error of the BEM method. In all numerical experiments, the sphere was discretized using a mesh with almost regular triangles based on a discretization of a projected subdivided icosaeder. Two meshes with $N=1280$ and $N=5120$ elements were used (see also Fig. 2), the radii of the sphere were set to $r=1$ and $r=2$, and the linear systems for all benchmark problems will be solved using an LU-decomposition implemented in LAPACK [3].
3.1.1 Unit Sphere
In the first experiments, the condition numbers and the relative errors of the BEM for different BM factors will be investigated for the example of plane wave scattering from the unit sphere that was discretized with $N=1280$ and $N=5120$ triangular elements (see also Fig. 2). The frequencies of interest were between $10$ Hz and $500$ Hz in steps of 10 Hz. For a frequency of 500 Hz, the average edge length of the elements of the mesh was slightly over the limit recommended by the 6-to-8 elements per wavelength rule of thumb [13]. Around the critical frequencies, which are given by the zeros of the spherical bessel function $j_{n}(k)$ and which are summarized in Tab. 1, 9 steps with step size $0.05$ Hz were made. Because of the discretization of the sphere’s surface and the fact that the solution is calculated at the midpoints of the elements, the critical frequencies for the BEM calculation are not exactly at the zeros of the Bessel functions. These frequencies were determined by finding the frequency with a large condition number close to the theoretical critical frequency. These values are close to the theoretical frequencies for spheres where the radii are determined by the mean norm of the collocation nodes.

### v2

@@ Abstract
The Burton-Miller method is a widely used approach in acoustics to enhance the stability of the boundary element method for exterior Helmholtz problems at so-called critical frequencies. This method depends on a coupling parameter $\eta$ and it can be shown that as long as $\eta$ has an imaginary part different from 0, the boundary integral formulation for the Helmholtz equation has a unique solution at all frequencies. A popular choice for this parameter is $\eta=\frac{{\mathrm{i}}}{k}$, where $k$ is the wavenumber. It can be shown that this choice is quasi optimal, at least in the high frequency limit. However, especially in the low frequency region, where the critical frequencies are still sparsely distributed, different choices for this factor result in a smaller condition number and a smaller error of the solution. In this work, alternative choices for this factor are compared based on numerical experiments. Additionally, a way to enhance the Burton-Miller solution with $\eta=\frac{{\mathrm{i}}}{k}$ for a sound hard scatterer in the low frequency region by an additional step of a modified Richardson iteration is introduced.
@@ 1 Introduction
The boundary element method (BEM, [17, 19]) is an often used numerical method to solve exterior scattering problems in acoustics. It is a well known fact that although the exterior Helmholtz equation has a unique solution, there are certain critical frequencies where the boundary integral equation (BIE) used by the BEM has not [7, 20]. As the number of these critical frequencies below a given wavenumber $k$ is proportional to $k^{3}$ (cf. [3]), this non-uniqueness poses a major problem for the Helmholtz BEM.
@@
Eq. (2) will be called dBIE in the rest of the paper. $\phi({\bf x})$ denotes the velocity potential at a point ${\bf x}$ on a smooth part of the boundary $\Gamma$, $v({\bf x}):=\frac{\partial\phi}{\partial{\bf n}_{x}}\left({\bf x}\right):=\nabla\phi({\bf x})\cdot{\bf n}_{x}$ and
@@
where ${\bf n}_{y}$ is the normal vector to the boundary $\Gamma$ at a point ${\bf y}\in\Gamma$. In Eqs. (1) and (2) it is assumed that the boundary $\Gamma$ is sufficiently smooth.
@@
Different values for the BM factor $\eta$ and its sign have been discussed in the past (cf. [2, 3, 12, 16, 18, 22, 24]), the most commonly used factor is $\eta=\frac{{\mathrm{i}}}{k}$, which for high frequencies is almost optimal in terms of having a small condition number for problems on the unit circle or the unit sphere. Since scattering problems involving the sphere have an analytic solution based on spherical harmonics and the spherical Bessel and Hankel functions, it is possible to derive the eigenvalues and the condition number of the linear system for the combined integral equation as a function of $\eta$, see also Section 2. Based on this analysis, it can be shown that $\eta=\frac{{\mathrm{i}}}{k}$ is quasi-optimal (cf. [2]).
@@
$=\left(O\left(\frac{1}{r}\right)+\mathcal{O}(k)\right)G({\bf x},{\bf y}),$
@@
$=\left(O\left(\frac{1}{r^{2}}\right)+\mathcal{O}(k/r)+\mathcal{O}(k^{2})\right)G({\bf x},{\bf y}),$
where $r=||{\bf x}-{\bf y}||$ (cf. [8, Eq. (15)]). The components of the dBIE contain a scaling with the wavenumber $k$, at least if $k$ is big enough so that the parts involving $\mathcal{O}(k)$ and $\mathcal{O}(k^{2})$ are bigger than the $\mathcal{O}\!\left(\frac{1}{r^{2}}\right)$ and $\mathcal{O}\!\left(\frac{1}{r}\right)$ parts. Using this viewpoint, multiplying the dBIE with $\frac{1}{k}$ balances the scale of BIE and dBIE. In 2D, a similar heuristic argument can be given when looking at the derivatives of the Green’s function $H_{0}(kr)$ with respect to $r$.
One can clearly see in Fig. 1 that for the example of plane wave scattering from a sound hard sphere, the Burton-Miller method has a positive effect on the stability of the solution of the scattering problem. However, in the low frequency region, the relative error between numerical and analytical solution away from the critical frequencies is, in general, smaller for the version without Burton-Miller. Using again the heuristic viewpoint, it seems that in the low frequency region, the $\mathcal{O}(k)$ factor becomes too small and the parts depending on the radius dominate the error.
The data in Fig. 1 was calculated for $N=1280$ almost regular triangles, using collocation with constant elements [13]. The highest investigated frequency of 500 Hz is slightly larger than suggested by the 6-to-8-elements per wavelength rule of thumb [15]. Still, the relative error at this frequency is acceptable.
@@
The example presented in Fig. 1 illustrates the problem of choosing a good BM-parameter if BEM solutions at low frequencies are needed. Examples where solutions for small wavenumbers are needed are the 2.5D BEM [10, 11] or the BEM for periodic structures using the Floquet transform [9]. The fact that for low frequencies the coupling factor $\eta=\frac{{\mathrm{i}}}{k}$ becomes very large, led to different approaches to define this factor for small $k$. One suggestion [2] is to use $\eta=\min(2,\frac{1}{k})\cdot{\mathrm{i}}$, which means for a sphere with radius $R=1$ m that above about 27 Hz the BM-factor is $\eta=\frac{{\mathrm{i}}}{k}$, below it is set constant. It was stated in [3] that “[…] 1 has often been used in practice”, where the authors referred to the imaginary part of the coupling factor. In [10], a factor $\eta=\frac{{\mathrm{i}}|k|}{\max(|k|^{2},50)}$ was suggested in connection with the 2D and 2.5D BEM. In [6], a BM factor of $\eta=\min(\frac{1}{3},\frac{\lambda}{D})\cdot{\mathrm{i}}=\min(\frac{1}{3},\frac{2\pi}{Dk})\cdot{\mathrm{i}}$ was used [^1] In fact, in [6] the authors scale the BIE, and the factor was adapted accordingly., where $D$ is the diameter of the scatterer. According to [6], this factor is based on optimizing the number of necessary GMRES iterations. [6] explicitly includes the diameter $D$ of the scatterer, whereas in some studies this scaling with the size of the object is implicitly assumed (cf. [2]: “If we denote $kR$ by $k^{*}$ and $\nu/R$ but $\nu^{*}$ then the problem is reduced to that of a unit sphere at the wave number $k^{*}$ with the coupling parameter $\nu^{*}$, a similar claim can be found in [12].).
@@
In this study, different choices of the Burton-Miller factor and their effect on the solution of a scattering problem will be investigated. As a benchmark example, the scattering of a plane wave on the unit sphere will be used. As there is some discussion about whether to adapt the BM factor to the size of the scatterer, one numerical experiment will look at the sphere with radius $R=2$ m. Also, a new approach based on a single step of a Modified Richardson iteration is proposed to enhance the solution of the Burton-Miller BEM in the low frequency region for sound hard problems.
In Section 2 an eigenvalue analysis of the BIE operators on the sphere is derived that helps in explaining some of the effects seen in Section 4. After a short motivation for the additional iteration step in Section 3, numerical benchmark tests are done in Section 4, where the solutions for different Burton Miller factors are compared with each other. The paper is finished with a short discussion in Section 5.
2 Eigenvalue-Analysis for the sphere
Following the approach proposed in [2], it is possible to use the orthogonality of the Spherical Harmonics to derive the eigenvalues of the different integral operators involved with the Helmholtz equation on a sphere with radius $R=1$ m as function of the wavenumber $k$:
$\lambda_{n}(G)$ $={\mathrm{i}}kh_{n}(k)j_{n}(k)\xrightarrow[k\rightarrow 0]{}\frac{1}{2n+1},$ (4)
$\lambda_{n}(H)=\lambda_{n}(H^{\prime})$ $=\frac{1}{2}+{\mathrm{i}}k^{2}h^{\prime}_{n}(k)j_{n}(k)\xrightarrow[k\rightarrow 0]{}-\frac{1}{2(2n+1)},$ (5)
$\lambda_{n}(E)$ $={\mathrm{i}}k^{3}h^{\prime}_{n}(k)j^{\prime}_{n}(k)\xrightarrow[k\rightarrow 0]{}-\frac{n(n+1)}{2n+1},$ (6)
where $j_{n},h_{n},j^{\prime}_{n},$ and $h^{\prime}_{n}$ are the spherical Bessel function of order $n$, the spherical Hankel function of the first kind, and their respective derivatives. From Eq. (5) it is easy to see that the hypersingular operator becomes unstable at low frequencies because one of the eigenvalues of this operator converges towards 0. A similar instability can be shown for the operator $\frac{1}{2}+H^{\prime}$. Both operators are involved in the dBIE for sound hard and sound soft problems, which indicates that for small wavenumbers the dBIE for both problems is unstable. It is also easy to show that the eigenvalues for the operators $G$ and $\frac{1}{2}-H$ are bounded and sufficiently away from 0, thus, the BIE for sound hard and sound soft problems stays stable for small wavenumbers and mesh sizes $N$ that are not extensively large. [^2] From operator theory it is known that for a linear operator $A:X\rightarrow Y$ the problem $A\phi=f$ becomes ill posed if $X$ has infinite dimension, which could negatively affect the BIE for the sound soft problem. However, for practical applications the mesh sizes are reasonable, so the ill-posedness is rather an academic problem.
Using the asymptotic expansions [1, 2] of the Bessel and Neumann functions for small arguments $(k\rightarrow 0)$
$j_{n}(k)$ $\approx\frac{k^{n}}{1\cdot 3\cdots(2n+1)}=\mathcal{O}(k^{n})$
$y_{n}(k)$ $\approx-\frac{1\cdot 3\cdots(2n-1)}{k^{n+1}}=\mathcal{O}(k^{-(n+1)})$
it is possible to investigate the behavior of the different solutions in the low frequency region. If one looks at the eigenvalues of the operators involved in the Burton Miller approach for the Dirichlet and the Neumann problems [^3] The values can be either found in [2] or directly derived by looking the values for the boundary integral operators and the fact that the Wronskian $j_{n}(k)h^{\prime}_{n}(k)-j^{\prime}_{n}(k)h_{n}(k)=\frac{i}{k^{2}}$.
$\lambda_{n}(D)$ $={\mathrm{i}}kh_{n}(k)\left(j_{n}(k)-\eta kj^{\prime}_{n}(k)\right),$
$\lambda_{n}(N)$ $={\mathrm{i}}k^{2}h^{\prime}_{n}(k)\left(j_{n}(k)+\eta kj^{\prime}_{n}\right).$
it becomes apparent that for small $k$ and $\eta=\frac{{\mathrm{i}}}{k}$ the eigenvectors of the operators diverge:
$\lambda_{n}(D)$ $={\mathrm{i}}k\left(\mathcal{O}(k^{n})+\mathcal{O}(k^{-(n+1)})\right)\left(\mathcal{O}(k^{n})+\mathcal{O}(k^{n-1})\right)=\mathcal{O}(k^{2n+1})+\mathcal{O}(1)+\mathcal{O}(k^{2n})+\mathcal{O}(k^{-1})$
$\lambda_{n}(N)$ $={\mathrm{i}}k^{2}\left(\mathcal{O}(k^{n-1})+\mathcal{O}(k^{-(n+2)})\right)\left(\mathcal{O}(k^{n})+\mathcal{O}(k^{n-1})\right)=\mathcal{O}(k^{2n+1})+\mathcal{O}(1)+\mathcal{O}(k^{2n})+\mathcal{O}(k^{-1})$ (7)
This explains the rise in the Burton-Miller solution in Section 4 towards $k=0$ if $\eta=\frac{{\mathrm{i}}}{k}$. If the Burton-Miller factor scales with $k$ or is kept constant in the low frequency region, this rise in the eigenvalues can be avoided.
3 Iterative Correction of the BM Solution for Sound Hard Problems
@@
It was illustrated in the previous section that in the low frequency region away from the critical frequencies, the Burton-Miller method produces larger errors than the solution of the original BIE. An alternative to using different Burton-Miller factors is presented in the following. As the operators and integral kernels involved with the dBIE are, in general, numerically more challenging than the operators involved with the BIE, it is plausible to assume that the error of a solution based purely on the BIE is smaller than the error for the Burton-Miller solution as long as the frequency is sufficiently far away from the set of critical frequencies, where the BIE has no unique solution (see also Fig. 1). This becomes also clear, when the eigenvalues for the different operators on the sphere are compared, see Section 2. For the BIE the eigenvalues of $\frac{1}{2}-H$ converge towards $\frac{n+1}{2n+1}$ for $k\rightarrow 0$ thus they stay away from zero. For the hypersingular operator and therefore, for the dBIE, Eq. (6) shows that for $n=0$ the operator becomes singular.
@@
In [21, Theorem 3] it is shown that if a system $\phi-K\phi=0$ with compact operator $K$ has only the trivial solution $\phi=0$, and if $\phi_{0}$ is a Galerkin-Petrov solution of $\phi=f+K\phi$, then $\phi_{1}=f+K\phi_{0}$ is a better approximation of the real solution provided that the number of ansatz functions $n$ is big enough. As the theorem is based on orthogonal projections on the ansatz and test function spaces, it cannot be applied to the collocation method directly. Also, the fact that the system needs to have a unique solution is not fulfilled at the critical frequencies. Nevertheless, this theorem serves as a motivation for using an additional iteration step in connection with the problem of solving wave scattering around a sound hard sphere:
@@
Note that the approach described in [21] is fitted to equations of the type $\phi=f+A\phi$, where $A$ is some linear operator. This is also a restriction of the approach proposed here that prohibits an efficient solution of a sound soft problem by this approach. Theoretically, it is possible to use the dBIE
$\frac{{\bf v}}{2}+{\bf H}^{\prime}{\bf v}={\bf v}_{\text{inc}}$
to construct a similar iteration
${\bf v}_{n+1}=2\left({\bf v}_{\text{inc}}-{\bf H}^{\prime}{\bf v}_{n}\right),$
however, the discussion in Section 2 with respect to the eigenvalues of the different boundary integral operators shows that for the example of the sphere one of the eigenvalues for the associated double layer potential $H^{\prime}$ is $-1/2$, see Eq. (5). Thus the stability of the dBIE deteriorates for small $k$ and an additional iteration step just based on this equation will not lower the error.
@@ 4.1 Sound Hard Sphere
As the first benchmark example, the standard problem of plane wave scattered at a sound hard sphere is used. This example has the advantage, that an analytic solution of the problem exists [23, Eq. (6.186)], and it is therefore possible to easily calculate the numerical error of the BEM method. In all numerical experiments, the sphere was discretized using a mesh with almost regular triangles based on a discretization of a projected subdivided icosaeder. Two meshes with $N=1280$ and $N=5120$ elements were used (see also Fig. 2), the radii of the sphere were set to $R=1$ m and $R=2\,$ m, and the linear systems for all benchmark problems will be solved using an LU-decomposition implemented in LAPACK [4]. In all computations it is assumed that the speed of sound and the density are given by $c=340$ m s-1 and $\rho=1.3$ kg m-3.
4.1.1 Sphere with radius $R=1$ m
In the first experiments, the condition numbers and the relative errors of the BEM for different BM factors will be investigated for the example of plane wave scattering from a sphere with radius $R=1$ m that was discretized with $N=1280$ and $N=5120$ triangular elements (see also Fig. 2). The frequencies of interest were between $10$ Hz and $500$ Hz in steps of 10 Hz. For a frequency of 500 Hz, the average edge length of the elements of the mesh was slightly over the limit recommended by the 6-to-8 elements per wavelength rule of thumb [15]. Around the critical frequencies, which are given by the zeros of the spherical Bessel function $j_{n}(k)$ and which are summarized in Tab. 1, 9 steps with step size $0.05$ Hz were made. Because of the discretization of the sphere’s surface and the fact that the solution is calculated at the midpoints of the elements, the critical frequencies for the BEM calculation are not exactly at the zeros of the Bessel functions. These frequencies were determined by finding the frequency with a large condition number close to the theoretical critical frequency. These values are close to the theoretical frequencies for spheres where the radii are determined by the mean norm of the collocation nodes.